# SPA: Towards A Computational Friendly Cloud-Base and On-Devices Collaboration Seq2seq Personalized Generation

### v1

@@ Abstract
Large language models(LLMs) have shown its outperforming ability on various tasks and question answering. However, LLMs require high computation cost and large memory cost. At the same time, LLMs may cause privacy leakage when training or prediction procedure contains sensitive information. In this paper, we propose SPA(Side Plugin Adaption), a lightweight architecture for fast on-devices inference and privacy retaining on the constraints of strict on-devices computation and memory constraints. Compared with other on-devices seq2seq generation, SPA could make a fast and stable inference on low-resource constraints, allowing it to obtain cost effiency. Our method establish an interaction between a pretrained LLMs on-cloud and additive parameters on-devices, which could provide the knowledge on both pretrained LLMs and private personal feature.Further more, SPA provides a framework to keep feature-base parameters on private guaranteed but low computational devices while leave the parameters containing general information on the high computational devices. [^1] Our code is public at https://github.com/OceannTwT/SPA
@@ 1 Introduction
Large Language Models (LLMs)(Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023) have demonstrated incredibly powerful capabilities in supporting downstream tasks. These models perform exceptionally well across various downstream tasks. Additionally, different LLMs such as GPT-4(OpenAI, 2023), Llama(Touvron et al., 2023), and ChatGLM(Zeng et al., 2022; Du et al., 2022) exhibit similar abilities in handling tasks across multiple dimensions(Bang et al., 2023). However, deploying and computing LLMs on resource-constrained devices introduces significant memory and computational pressure on these devices(Alizadeh et al., 2023). Therefore, researching reliable methods for deploying LLMs on-devices becomes critically urgent.
@@
To address the issue of model deployment on-devices, several methods have propose different approachs to adapt LLMs on edge devices. Subformer(Reid et al., 2021a) reduces the parameter sizes the model needs to store and the computational FLOPS required by sharing parameters. Similarly, EdgeFormer(Ge et al., 2022) further explores the impact of different architectures on model deployment on-devices and proposes a balanced sharing scheme. LLMCad(Xu et al., 2023) utilizes a collaborative model configuration mode to achieve lightweight model deployment, where a smaller LLM generates the simplest words, while a high-precision LLM verifies these words and corrects errors.
Furthermore, prior researches provide some useful improvements in parameter efficient fine-tuning LLMs(Houlsby et al., ; Hu et al., 2021; Lee et al., 2022). These methods efficiently utilize the additional parameter space to adapt LLMs to downstream tasks flexibly while retaining the inherent capabilities of the models. Notably, efficient fine-tuning techniques for large language models are based on adapter strategies, where adapters integrate with the base model to produce effective capabilities for the required downstream tasks. This suggests that in fine-tuning large language models, adapters can be separated from the core model, enabling storage separation of these components(Hu et al., 2023). This separation brings significant advantages in device storage, as only a small portion of the original model parameters needs to be stored, facilitating the retention of adapters tailored for specific downstream tasks.
Considering the characteristic of fewer adapter parameters, deployment on-devices can use adapters as incremental deployments(Lin et al., 2020), with adapters serving as learned components for downstream task. It integrates incremental knowledge into the architecture of the existing model to obtain inherent characteristics of users. Leveraging the additional adapters, effective integration between this incremental knowledge and the base large language model can strongly support downstream tasks and encapsulate the process of obtaining user attributes. However, this design aspect needs to consider the requirements of inter-device data transmission(Ren et al., 2019). When deploying adapters on-device, complexities and inference delays related to inter-device communication arise due to the isolation of the language model and its adapters on different devices. Addressing how to reduce the latency of transferring this portion of data presents a major challenge for designs of separated models.
@@ 2.1 Parameter-Efficient Transfer Learning
In recent years, efficient parameter-efficient transfer learning has become a popular method for fine-tuning large language models. It achieves performance close to full fine-tuning by refining a subset of parameters. Among these methods, adapters, as introduced by (Houlsby et al., 2019), incorporate two additional feed-forward layers, enhancing their adaptability to downstream tasks. Similarly, LoRA (Hu et al., 2021) employs a rank reduction followed by rank promotion method on the original language model. In the context of language model fine-tuning, prefix tuning (Li and Liang, 2021) involves introducing a trainable prefix vector before the keys and values in the multi-head attention mechanism, producing effects similar to the first two methods.
The MAM adapter (He et al., 2021) synthesizes these techniques and identifies superior results in parameter fine-tuning. Moreover, LST (Sung et al., 2022) provides a method for saving space during parameter fine-tuning. It achieves optimal parameter fine-tuning with the same memory consumption by leveraging adapters on the device. This approach significantly reduces memory usage when training large language models on edge devices. Meanwhile, the QLora (Dettmers et al., 2023) method combines quantization rules with LoRA’s rank promotion and demotion scheme, optimizing training resources and space savings.
Furthermore, the CuQA (Lee et al., 2022) method offers a classifier that selects between the original language model and edge models based on the relevance of training and testing data during training. Through this selection process, we observe that efficient parameter fine-tuning methods can effectively embed new knowledge, thereby expanding existing language models to incorporate new knowledge bases in support of specific tasks. Our method leverages the parameter-saving capabilities of efficient parameter-efficient transfer learning and the adaptive information contained in parameter fine-tuning inspired by CuQA. We provide robust support for various generative tasks, demonstrating enhanced capabilities.
@@ 2.2 On-device ML optimization
For on-device sequence-to-sequence models, various techniques have been adopted to reduce the parameters of pre-trained models, including methods such as quantization(Liu et al., 2018; Li et al., 2023) to reduce model size, knowledge distillation(Rao et al., 2023; Wang and Yoon, 2021) techniques that learn from large models and transfer knowledge to smaller models, and techniques such as compression(Frantar et al., 2022; Pope et al., 2023) and pruning(Bolya et al., 2022; Cai et al., 2022; Kim et al., 2022) to reduce the number of model parameters. Some methods like Edgeformer (Reid et al., 2021b) achieve model reduction by sharing model parameters. Low-rank decomposition is also used to reduce model size (Cahyawijaya, 2021). Methods like Delight (Mehta et al., 2020) achieve model size reduction by compressing attention mechanisms. These techniques are adjusted and adapted based on the parameters of the Transformer (Vaswani et al., 2017) architecture, and they yield effective parameter reduction results for models with relatively fewer parameters.
@@
Figure 2: Cloud-base and on-devices collaboration model: SPA.
@@ 3 Side Plugin Adaption
Given the computational and storage capabilities on the edge, and the need to ensure strong data privacy during model inference, our concerns can be broken down into the following points:
@@
  • Is there a possibility of information leakage between the edge and cloud models during their interactions?
@@
These are some fundamental requirements and abilities for a small on-device model that guarantees data privacy. Considering the need to learn user features and context-related information for text generation, we propose adding a classifier after the output of the transformer. We explore two settings for this classifier, comparing them to the original edge adapter’s results to assess downstream task support.
@@
Here we keep $f(x)$ frozen and train the side model $g(x)$ and classifier $\sigma$. Unlike a fixed classifier $\sigma^{\prime}$ for text similarity on the original model, our linear-layer-based classifier $\sigma$ can determine if a predicted word exists in the knowledge base of the base language model. This approach reduces inference times for highly repetitive word sequences, storing more general information in the cloud and generating user-specific features on the edge device.
From a privacy view, this classifier ensures that the cloud can only access the current classifier binary choice information, not the edge model’s parameters. Thus, during inference, the cloud only receives a signal to involve the edge when necessary, significantly enhancing personal data privacy. Compared to EdgeFormer, our model’s main bottleneck lies in the communication latency (Ren et al., 2019) between the edge and server. This can be managed by controlling communication frequency and times for an optimal value.
EdgeFormer (Reid et al., 2021b) scales down the original model by factors like four or eight, sharing parameters to achieve a smaller model with similar capabilities and fewer FLOPs. However, this design demands greater computational and proficiency requirements for edge devices. While our methods set the edge model small, its prediction speed relative to total latency is quite small. In this case, the edge devices computational ability would not be a bottleneck for predicting the results on-devices.
For communication between cloud servers and edge models, we adopt the architecture used in LST. LST improves memory capacity from back propagate. For prediction, if our transformer has $L$ layers, Parallel adapter (He et al., 2021) or LoRA methods would cause a large cost on communication latency. Specifically, the latency including three parts. $T_{pretrained}$ indicates the original pretrained model computing latency. $T_{on-devices}$ indicates the edge devices model computing latency, it could be Adapter part, LoRA part or LST part for on-devices inference. What’s more, $T_{net}$ is the network latency when transmitting data between cloud and devices. Then, the total latency could be evaluated as:
@@
$T_{net}=M\times(\tau+T_{data})$ (6)
Where M is the transmission times for one token prediction, it would be changed for different architecture, because we have to transfer back the hidden state from side model to base model. LST effectively merges the process in scenarios where edge-cloud connectivity is involved.
@@
Table 1: The Rouge-L scores on LST model, one-shot instruct and SPA model on various datasets
@@
4 Experiments Setup
4.1 Datasets
We evaluated our approach on the following training dataset.
  1. 1.
    XSum (Narayan et al., 2018): There are 227,000 summaries of online articles generated by the BBC available on XSum, and we assessed the relationships between the articles and their corresponding summaries.
  2. 2.
    CNN Daily Mail (Chopra et al., 2016): The CNN Daily Mail dataset consists of 312,000 summaries of CNN Daily Mail articles and its origin articles.
  3. 3.
    CoQA (Reddy et al., 2019): CoQA datasets contains 127,000 questions encompassing 8,000 conversation-based question-answer pairs, constituting a large-scale QA system spanning various domains.
  4. 4.
    SciQ (Welbl et al., 2017): The SciQ dataset comprises 13,679 publicly available scientific exam questions covering diverse disciplines such as biology, chemistry, physics, and others. Answers to specific questions are obtained through contextual information.
@@ 4.1 Datasets
Model % Param M Diff rates Diff Pick rates
LLaMA-7B + SPA 0.21 0.62 78.2 56.8
Table 4: The Side Model Effect evaluation.
4.2 Implement Details
[…]
5 Experiments
5.1 Latency on Prediction
@@ 5.3 Analysis on Training Data Sizes
Data sizes Xsum CNNDM CoQA SciQ
$\text{SPA}_{\text{small}}$ 10000 20000 4000 5000
$\text{SPA}_{\text{medium}}$ 100000 150000 40000 10000
$\text{SPA}_{\text{full}}$ 227000 312000 127000 13679
Table 5: Different training data sizes on datasets.
@@
Usage Percentage 82.3 71.5 61.8
Table 7: Evaluations on side model ratios.
@@
55551010101015151515202020202424242428282828323232323636363640404040epochesRouge-LXSumLSTParallel LST
55551010101015151515202020202525252529292929333333333737373741414141epochesRouge-LCNN-DMLSTParallel LST
@@ 5.4 Different Setting on Side Plugin
The final setup involves the existing LST model, which can be viewed as relying entirely on the edge adapter for result generation.
By comparing our model with these three setups, we aim to determine which configuration best aligns with the requirement for consistency between efficiency and capability that the model needs to achieve.The capabilities demonstrated by different experiments are summarized in Table 3.
The model’s enhancement in downstream tasks primarily relies on the adjustments made by the adapter to capture the incremental differences. As illustrate on Table 3, While the introduction of the classifier in the SPA model leads to a slight performance drop compared to the scenario where adapters are integrated in all cases, the performance reduction is relatively minimal. However, the improvement in inference speed is substantial. Considering our requirement for quick responses at the edge, such a classifier aligns better with our vision and needs. Simultaneously, for more fundamental information, predictions are made using the original model. The classifier significantly reduces the latency caused by continuous calls to the edge-side model during predictions, ensuring that the model learns both user and task features effectively.
We evaluate the performance of the sequence to sequence task between the original model and the model after parallel processing, and make corresponding comparisons under different datasets, and determine its performance under the corresponding task by the Rouge-L parameter.
@@ 6 Conclusions
7 Limitation
Our model is an cloud-based and on-devices cooperative model, so it embodies a fairly strong inference ability on some downstream tasks, while relying on the high-parameter nature of the large model, the end-to-end prediction can also have a very good result. Our model still has some shortcomings. For example, in such a collaborative process, its delay is still unavoidable. In some scenarios that require high network resources, our model may still not be worked for reasoning and training in such scenarios.
In addition, there may be some privacy issues in our scenarios, because our corpus is ultimately delivered to the cloud, so the cloud services can still see the corresponding corpus content. However, we still protect the privacy of the model to a certain extent, because the parameters related to the model are kept on our local side, and we only need these parameters to realize the inference of the model for the user features. These problems also deserve to be further investigated and addressed in future work.
Ethics Statement
Our approach utilizes open-source datasets for training and evaluating the model. All data are openly accessible. We leverage open-source code and weights for Llama. Due to the hallucination issue of large language models, some of the generated content may contain factual errors and reasoning errors. SPA offers a potential on-devices LLM solution based on cloud service and edge model collaboration. Our work strictly adheres to the license and policies of released LLMs and publicly available datasets.

### v3

@@ Abstract
Large language models(LLMs) have shown its outperforming ability on various tasks and question answering. However, LLMs require substantial memory storage on low-resource devices. More critically, the computational speed on these devices is also severely limited. In this paper, we propose SPA(Side Plugin Adaption), a lightweight architecture for fast on-devices inference on the constraints of strict on-devices computation and memory constraints. Compared with other on-devices seq2seq generation, SPA could make a fast and stable inference on low-resource constraints, allowing it to obtain cost effiency. Our method establish an interaction between a pretrained LLMs on-cloud and additive parameters on-devices, which could provide the knowledge on both pretrained LLMs and featured personal feature. Further more, SPA provides a framework to keep feature-base parameters on low computational devices while leave the parameters containing general information on the high computational devices.  [^1] Our code is public at https://github.com/OceannTwT/spa
Keywords:
Personalized LLM Cloud-device Collaboration Inference Acceleration.
@@ 1 Introduction
Large Language Models (LLMs)[4, 8, 33] have demonstrated incredibly powerful capabilities in supporting downstream tasks. These models perform exceptionally well across various downstream tasks. Additionally, different LLMs such as GPT-4[25], Llama[33], and ChatGLM[10, 38] exhibit similar abilities in handling tasks across multiple dimensions[2]. However, deploying and computing LLMs on resource-constrained devices introduces significant memory and computational pressure on these devices[1]. Therefore, researching reliable methods for deploying LLMs on-devices becomes critically urgent.
@@
To address the issue of deploying large language models (LLMs) on devices, various methods have been proposed to adapt them for edge devices. Subformer [29] reduces storage and computation requirements by sharing parameters. Similarly, EdgeFormer [12] proposes a balanced sharing scheme, exploring the impact of different architectures on device deployment. LLMCad [37] employs a collaborative model configuration to achieve lightweight deployment, where a small model generates simple words and a high-accuracy model verifies and corrects errors.
Additionally, research has proposed parameter-efficient fine-tuning methods for LLMs [14, 15, 18], utilizing adapter strategies to flexibly adapt to downstream tasks while preserving model capabilities. Adapters are stored separately from the base model and require only a few parameters to customize for downstream tasks [16]. For on-device deployment, adapters can be used for incremental deployment, integrating incremental knowledge into existing models to support downstream tasks and obtain user attributes [21]. However, this approach requires consideration of data transmission requirements between devices [31], and reducing transmission latency is a significant challenge in the design of separated models.
@@ 2.1 Parameter-Efficient Transfer Learning
In recent years, efficient parameter-efficient transfer learning has become a popular method for fine-tuning large language models. It achieves performance close to full fine-tuning by refining a subset of parameters. Among these methods, adapters, as introduced by [14], incorporate two additional feed-forward layers, enhancing their adaptability to downstream tasks. Similarly, LoRA [15] employs a rank reduction followed by rank promotion method on the original language model. In the context of language model fine-tuning, prefix tuning [19] involves introducing a trainable prefix vector before the keys and values in the multi-head attention mechanism, producing effects similar to the first two methods.
The MAM adapter [13] synthesizes these techniques and identifies superior results in parameter fine-tuning. Moreover, LST [32] provides a method for saving space during parameter fine-tuning. It achieves optimal parameter fine-tuning with the same memory consumption by leveraging adapters on the device. This approach significantly reduces memory usage when training large language models on edge devices. Meanwhile, the QLora[9] method combines quantization rules with LoRA’s rank promotion and demotion scheme, optimizing training resources and space savings.
Furthermore, the CuQA[18] method offers a classifier that selects between the original language model and edge models based on the relevance of training and testing data during training. Through this selection process, we observe that efficient parameter fine-tuning methods can effectively embed new knowledge, thereby expanding existing language models to incorporate new knowledge bases in support of specific tasks. Our method leverages the parameter-saving capabilities of efficient parameter-efficient transfer learning and the adaptive information contained in parameter fine-tuning inspired by CuQA. We provide robust support for various generative tasks, demonstrating enhanced capabilities.
@@ 2.2 On-device ML optimization
For on-device sequence-to-sequence models, various techniques have been adopted to reduce the parameters of pre-trained models, including methods such as quantization[20, 22] to reduce model size, knowledge distillation[27, 35] techniques that learn from large models and transfer knowledge to smaller models, and techniques such as compression[11, 26] and pruning[3, 6, 17] to reduce the number of model parameters. Some methods like Edgeformer [30] achieve model reduction by sharing model parameters. Low-rank decomposition is also used to reduce model size [5]. Methods like Delight [23] achieve model size reduction by compressing attention mechanisms. These techniques are adjusted and adapted based on the parameters of the Transformer [34] architecture, and they yield effective parameter reduction results for models with relatively fewer parameters.
@@ 3 Side Plugin Adaption
Given the computational and storage capabilities on the edge, our concerns can be broken down into the following points:
@@
Figure 2: Cloud-base and on-devices collaboration model: SPA.
These are some fundamental requirements and abilities for a small on-device model. Considering the need to learn user features and context-related information for text generation, we propose adding a classifier after the output of the transformer. We explore two settings for this classifier, comparing them to the original edge adapter’s results to assess downstream task support.
@@
Here we keep $f(x)$ frozen and train the side model $g(x)$ and classifier $\sigma$. Unlike a fixed classifier $\sigma^{\prime}$ for text similarity on the original model, our linear-layer-based classifier $\sigma$ can determine if a predicted word exists in the knowledge base of the base language model. This approach reduces inference times for highly repetitive word sequences, storing more general information in the cloud and generating user-specific features on the edge device. Compared to EdgeFormer, our model’s main bottleneck lies in the communication latency[31] between the edge and server. This can be managed by controlling communication frequency and times for an optimal value.
EdgeFormer[30] scales down the original model by factors like four or eight, sharing parameters to achieve a smaller model with similar capabilities and fewer FLOPs. However, this design demands greater computational and proficiency requirements for edge devices. While our methods set the edge model small, its prediction speed relative to total latency is quite small. In this case, the edge devices computational ability would not be a bottleneck for predicting the results on-devices.
For communication between cloud servers and edge models, we adopt the architecture used in LST. LST improves memory capacity from back propagate. For prediction, if our transformer has $L$ layers, Parallel adapter[13] or LoRA methods would cause a large cost on communication latency. Specifically, the latency including three parts. $T_{pretrained}$ indicates the original pretrained model computing latency. $T_{on-devices}$ indicates the edge devices model computing latency, it could be Adapter part, LoRA part or LST part for on-devices inference. What’s more, $T_{net}$ is the network latency when transmitting data between cloud and devices. Then, the total latency could be evaluated as:
@@
$T_{net}=\mathcal{M}\times(\tau+T_{data})$ (6)
Where $\mathcal{M}$ is the transmission times for one token prediction, it would be changed for different architecture, because we have to transfer back the hidden state from side model to base model. LST effectively merges the process in scenarios where edge-cloud connectivity is involved.
4 Experiments Setup
4.1 Datasets
We evaluated our approach on the following training dataset.
  1. 1.
    XSum[24]: There are 227, 000 summaries of online articles generated by the BBC available on XSum, and we assessed the relationships between the articles and their corresponding summaries.
  2. 2.
    CNN Daily Mail[7]: The CNN Daily Mail dataset consists of 312, 000 summaries of CNN Daily Mail articles and its origin articles.
  3. 3.
    CoQA[28]: CoQA datasets contains 127, 000 questions encompassing 8, 000 conversation-based question-answer pairs, constituting a large-scale QA system spanning various domains.
  4. 4.
    SciQ[36]: The SciQ dataset comprises 13, 679 publicly available scientific exam questions covering diverse disciplines such as biology, chemistry, physics, and others. Answers to specific questions are obtained through contextual information.
4.2 Implement Details
[…]
5 Experiments
Table 1: The Rouge-L scores on LST model, one-shot instruct and SPA model on various datasets.
@@ 5 Experiments
5.1 Latency on Prediction
@@
Table 4: Different training data sizes on datasets.
Data sizes Xsum CNNDM CoQA SciQ
$\text{SPA}_{\text{small}}$ 10000 20000 4000 5000
$\text{SPA}_{\text{medium}}$ 100000 150000 40000 10000
$\text{SPA}_{\text{full}}$ 227000 312000 127000 13679
@@ 5.3 Analysis on Training Data Sizes
Table 6: Evaluations on side model ratios.
Usage Percentage 82.3 71.5 61.8
@@
55551010101015151515202020202424242428282828323232323636363640404040Num of TrainningRouge-LXSumLSTParallel LST
55551010101015151515202020202525252529292929333333333737373741414141Num of TrainningRouge-LCNN-DMLSTParallel LST
@@ 5.4 Different Setting on Side Plugin
Table 7: The Side Model Effect evaluation.
Model % Param M Diff rates Diff Pick rates
LLaMA-7B + SPA 0.21 0.62 78.2 56.8
@@
The final setup involves the existing LST model, which can be viewed as relying entirely on the edge adapter for result generation. By comparing our model with these three setups, we aim to determine which configuration best aligns with the requirement for consistency between efficiency and capability that the model needs to achieve. The capabilities demonstrated by different experiments are summarized in Table 3.
The improvement of model performance in downstream tasks primarily relies on the adjustments made by the adapter to capture the incremental differences. As illustrate on Table 3, While the introduction of the classifier in the SPA model leads to a slight performance drop compared to the scenario where adapters are integrated in all cases, the performance reduction is relatively minimal. However, the improvement in inference speed is substantial. Considering our requirement for quick responses at the edge, such a classifier aligns better with our vision and needs. Simultaneously, for more fundamental information, predictions are made using the original model. The classifier significantly reduces the latency caused by continuous calls to the edge-side model during predictions, ensuring that the model learns both user and task features effectively.
We evaluate the performance of the sequence to sequence task between the original model and the model after parallel processing, and make corresponding comparisons under different datasets, and determine its performance under the corresponding task by the Rouge-L metrics.